The combined SAT scores for the students at a local high school are normally distributed with a mean of 1450 and a standard deviation of 302. The local college includes a minimum score of 2265 in its admission requirements. What percentage of students from this school earn scores that satisfy the admission requirement? P(X> 2265) Enter your answer as a percent accurate to 1 decimal place (do not enter the % sign). Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted. Get help: Video Video License Points possible: 5 This is attempt 4 of S. Score on last attempt: 0. Score in gradebook: 0 Submit

the combined SAT scores for the students at a local high school are normally distributed with a mean of 1466
Answer #1

\mu = 1450

\sigma = 302

To find P(X>2265):
Z = (2265 - 1450)/302 = 2.6987

Table of Area Under Standard Normal Curve gives area = 0.4965

So,

P(X>2265) = 0.5 - 0.4965 = 0.0035= 0.4 %

So,

Answer is:

0.4